MOA-II Obsevation Log Monthly Summary 2007/07
Return to index
Previous month
Next month
TARGET | RUN | TOTAL RUN | TOTAL FIELD | TOTAL EXPOSURE (sec) |
GB | 8060-9411 | 1352 | 5516 | 311465 |
LMC | 1581-1609 | 29 | 280 | 72150 |
SMC | 393-414 | 22 | 42 | 10400 |
ALERT | 358-418 | 61 | 61 | 3660 |
GRB | - | 0 | 0 | 0 |
TOTAL | --- | 1464 | 5899 | 397675 (110h00m1675s) |
GB FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
285 | 193 | 195 | 590 | 818 | 166 | 185 | 173 | 827 | 165 | 196 | 176 |
Field13 | Field14 | Field15 | Field16 | Field17 | Field18 | Field19 | Field20 | Field21 | Field22 | Field23 |
165 | 159 | 151 | 160 | 155 | 151 | 151 | 149 | 142 | 164 | 0 |
LMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
21 | 21 | 20 | 20 | 18 | 21 | 23 | 23 | 23 | 17 | 19 | 19 |
Field13 | Field14 | Field15 | Field16 |
17 | 18 | 0 | 0 |
SMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
22 | 20 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
SPECIAL TARGET NAME
Category | Target | Nrun |
ALERT | OGLE2007BLG363 | 14 |
ALERT | OGLE2007BLG373 | 47 |
CCD TEMPERATURE
Previous month
Next month

VACUUM
Previous month
Next month

TELESCOPE TEMPERATURE
Previous month
Next month

COMMENTS
2007/07/1: DARK D2415(40sec), D2416(30sec), D2417(150sec) are OK. We took the GB for 40sec and 30sec. / night time = 46500(sec)
2007/07/2: No observation. We took Flat images. Dark D2419 (15sec) is OK. / night time = 46500(sec)
2007/07/3: DARK D2422(60sec), D2423(200sec) are OK. / night time = 46440(sec)
2007/07/4: Dark D2426 (60sec) is OK. A new event posted in MOA275 was found by the real-time event check tools. / night time = 46380(sec)
2007/07/5: No observation. / night time = 46320(sec)
2007/07/6: Dark D2428(60sec), D2429(200sec) are OK. / night time = 46320(sec)
2007/07/7: DARK D2430(60sec), D2431(200sec), D2432(7sec), D2433(10sec), D2434(15sec), D2435(30sec) are OK. We observed gb4 for MOA233 quite high intensively, but it is possibly a normal event. / night time = 46260(sec)
2007/07/8: DARK D2437(300sec), D2438(60sec), D2439(10sec) are OK. / night time = 46260(sec)
2007/07/9: DARK D2443(60sec) is OK. / night time = 46140(sec)
2007/07/10: DARK D2447(60sec) is OK. A new vacuum sensor (IMR265) was possibly broken (at 20:15 NZST). Pumping. / night time = 46080(sec)
2007/07/11: DARK D2449(60sec), D2450(300sec) are OK. We observed a special field for OGLE363. / night time = 46080(sec)
2007/07/12: Dark D2451(60sec), D2452(300sec) are OK. An anomaly seemed to be found in MOA305 on gb1 and we observed this intensively, but it was wrong idea. / night time = 45960(sec)
2007/07/13: DARK D2453(300sec), D2454(60sec) are OK. We observed a special field for OGLE 373. We exchanged vacuum sensor. / night time = 45960(sec)
2007/07/14: DARK D2455(60sec), D2456(300sec) are OK. We observed gb4 quite intensively for MOA312. / night time = 45840(sec)
2007/07/15: DARK D2457(300sec), D2458(60sec) are OK. / night time = 45780(sec)
2007/07/16: DARK D2460(60sec), D2461(300sec) are OK. The filter does not move. / night time = 45720(sec)
2007/07/17: DARK D2463(60sec), D2464(300sec) are OK. / night time = 45660(sec)
2007/07/18: DARK D2468(60sec), D2470(300sec) are OK. / night time = 45540(sec)
2007/07/19: DARK D2471(60sec), D2472(300sec) are OK / night time = 45480(sec)
2007/07/20: DARK D2478(60sec), D2479(300sec) are OK. We observed gb1 as frequent as gb5,9 for OGLE399(anormaly). / night time = 45420(sec)
2007/07/21: DARK D2483(60sec), D2484(300sec) are OK. We observed only gb3 for MOA308(anormarly). We had electric power failure for 10 minutes. / night time = 45300(sec)
2007/07/22: DARK D2487 (60sec), D2488 (300sec) are OK. Intensive observation of gb20 for MOA253 is made (B9276-9280). / night time = 45240(sec)
2007/07/23: DARK D2489(60sec), D2490(300sec) are OK. / night time = 45120(sec)
2007/07/24: Windy and cloudy. / night time = 45060(sec)
2007/07/25: DARK D2494 (60sec), D2495 (300sec) / night time = 44940(sec)
2007/07/26: DARK D2497 (60sec), D2498 (30sec), no observation / night time = 44820(sec)
2007/07/27: Moon was at the center of bulge. Cloudy. DARK D2502 (60sec), D2503 (300sec) / night time = 44760(sec)
2007/07/28: DARK D2507 (60sec), D2508 (300sec), D2509 (150sec) are OK. / night time = 44640(sec)
2007/07/29: Bad weather. No observation. / night time = 44520(sec)
2007/07/30: DARK D2512, D2513, D2514 are OK. Coudy and bright moonlight. / night time = 44460(sec)
2007/07/31: Dark D2517(60sec), D2518(100sec), D2519(300sec) are OK. / night time = 44340(sec)
TOTAL NIGHT TIME: 1413840(sec)
OPERATING RATE: 44.8 %
GB: 37.6 %
LMC: 5.9 %
SMC: 0.9 %
ALERT: 0.4 %
GRB: 0.0 %
Return to index
Previous month
Next month